11. Perform the following hexadecimal additions and subtractions. Assume the numbers are stored in 32-bit 2’s complement binary numbers. Indicate the sign of the answer and whether overflow occurs.

a. BBCA270C + AE223464        

b. E3BA265F + E045B9A9        

c. E9B20F5D – FE605C8D         

d. 5FCA5243 – AE223464        
Answer #1

a)  BBCA270C + AE223464 = 169EC5B70
now the value of 169EC5B70
in 2's complement is :- 01 0110 1001 1110 1100 0101 1011 0111 0000 (34 bit req)

so overflow occur .

sign is +ve

b)  E3BA265F + E045B9A9 = 1C3FFE008

now the value of 1C3FFE008 in 2' complement

01 1100 0011 1111 1111 1110 0000 0000 1000 (34 bit req)

so overflow occur

c) E9B20F5D – FE605C8D         =  -14AE4D30

now the value in 2' complement

1110 1011 0101 0001 1011 0010 1101 0000

no overflow

d. 5FCA5243 – AE223464 = -4E57E221

now the value in 2' complement is

10110001101010000001110111011111

no overflow